RC meeting and Run Plan, Sunday February 5th
From SBS wiki
Return to the RC Meeting Calendar
Contents
RC Report
- https://logbooks.jlab.org/entry/4124877
- RC: Rachel Montgomery, rmontgom@jlab.org
Summary
- Day and Swing shifts yesterday were productive with good data taking running production on polhe3
- Owl shift beam was down for most of the time as they tried to bring Hall C back up.
- It looked like we were good to go with beam around the time the RC report went out, but now we are back to no beam and OPS are investigating BLM trips again.
Current Status
- Currently waiting on beam.
Run Plan
- Continue production running at 45uA and polHe3
Data Collection and Polarisaton Status
- 23.8C of uncorrected charge for this kinematic (26.1C of corrected charge)
- Current target polarisation 53.2% (no beam for a while). The polarisation has been really great lately during production running!
Other Topics/Status Updates
- MCC will do recovery day Tue next week and likely will accelerator tuning Thursday next week. I'm aware of two access requests right now, let me know of any others.
- Yesterday during swing we changed ps1 to 2 from -1. The event rate is now about 350Hz and live time is still ~100%. Sean updated the golden plots in the online display. HCal has some visible background now in the 2D ADC v TDC time colz plot, we checked with Jiwan who confirmed it makes sense.
Shift Schedule
- Shift Schedule Page
- Opening for SL on Fri 10th Feb swing, other than that looking great over the next week.
Requested Accesses to the Hall
- Looking to put SBS GEMs in DAQ during day shifts, need to fix a cable in Hall first
- Target team for oven heater check
Special Requests
- DC power did not access Hall during maintenance day for moeller magnet, would be good to coordinate in the future. I have asked the PD.
Notices for Shift Crew
- Check the How To page
- Pay attention to the beam restoration procedures in the How To
- there is a procedure if beam has been away for 30minutes, and another procedure if machine has been down for tuning or something which may change beam profile. Communicate with MCC to figure out why beam has been down and to figure out which kind of restoration you should do.
- SL:
- production runs (1-1.5hr) and replays (50k replay for each run, ~two runs also with 200k replays per shift if possible)
- TOs:
- NMR every 4 hours
- Note the instructions for performing an NMR if there has been any changes to magnetic fields (https://logbooks.jlab.org/entry/4124321)
- Contact RC any time for any questions